OpenStudy (anonymous):

|x + 7| = 6 help with absolute value

OpenStudy (mathstudent55):

When you have an absolute value equation of the type you have here, which is the absolute value of an expression equals a number do the following: 1. Write two equations separated by the word "or" 2. The first equation is simply the equation you were given but drop the absolute value signs. 3. The second equation is: take the negative of what is inside the absolute value signs, drop the absolute value signs and set that equal to zero. 4. Solve both equations and write both answers separated by word "or"

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

First equation is: x + 7 = 6 Then word "or" Second equation is -(x + 7) = 6 Put it all together and you have x + 7 = 6 or -(x + 7) = 6 Now solve both equations x = -1 or x + 7 = -6 x = -1 or x = -13

OpenStudy (mathstudent55):

You're welcome